# Communication-Efficient Federated Optimization over Semi-Decentralized Networks

### v2

@@ 5.1 Logistic regression with nonconvex regularization
To investigate communication efficiency of PISCO, we conduct experiments on logistic regression with a nonconvex regularization term [${}^{+}$19] using the a9a dataset [CL11]. Given the model parameter $\boldsymbol{x}$ and data sample $\boldsymbol{z}=(\boldsymbol{a},y)$, the empirical loss $\ell(\boldsymbol{x};\boldsymbol{z})$ is defined as:
@@
First, we study the influence of the agent-to-server communication probability $p$ on the training and test performance. To this end, we vary the probability $p$ from $\{1,1/10^{0.5},1/10^{0.75},1/10,1/10^{1.25},1/10^{1.5},1/10^{1.75},1/10^{2},0\}$ and presents the number of communication rounds of PISCO with different $p$ to achieve $0.05$ training accuracy (i.e., $\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}\leq 0.05$) and $80\%$ test accuracy ($\geq 95\%$ of the peak accuracy within 1000 communication rounds), in Figure 4.
@@
Figure 4: The number of agent-to-agent and agent-to-server communication rounds required to achieve $5\%$ training accuracy (the left panel) and $80\%$ test accuracy (the right panel) for PISCO with $T_{o}=1$ and different $p\in\{0,10^{-2},10^{-1.75},10^{-1.5},10^{-1.25},10^{-1},10^{-0.75},10^{-0.5},1\}$. Here, the blue (red) dotted line represents the number of agent-to-agent (agent-to-server) communication rounds that PISCO with $p=0$ (with $p=1$) requires.
@@
5.2 One-hidden-layer neural network training
@@
To verify the robustness of PISCO to diverse topologies, we consider a well-connected Erdős-Rényi topology with a connectivity probability of $0.3$ (corresponding to $\lambda_{w}=0.38$) and a disconnected Erdős-Rényi topology with a connectivity probability of $0.1$ (corresponding to $\lambda_{w}=0$). To simulate the highly data-heterogeneous scenario, we evenly split the sorted MNIST dataset to $n=10$ agents, where agent $i$ will receive the training data associated with label $i$, for $i\in[n]$. Moreover, we set the batch size $b=100$, the number of local updates $T_{o}=10$ and the agent-to-server communication probability $p\in\{1,1/\sqrt{n},1/n,0\}=\{1,10^{-0.5},10^{-1},0\}$. To reduce the impact of randomness, we run every experiment with 3 different seeds and show the average results in Figure 6.
(b) A disconnected Erdős-Rényi graph with the $0.1$ connectivity probability.
@@ 6 Conclusions
In this paper, we develop a communication-efficient algorithm called PISCO for solving federated nonconvex optimization over semi-decentralized networks, which enjoys the linear speedup of local updates and addresses data dissimilarity without any additional assumptions. By leveraging the heterogeneous communication model, PISCO largely reduces communication overheads in terms of the network dependency with a few agent-to-server communication rounds, particularly evident in poorly-connected networks. Both theoretical guarantees and empirical experiments underscore PISCO’s outstanding communication efficiency and robustness to data heterogeneity and various network topologies. In the future, it will be of interest to incorporate variance reduction [LLC22] and communication compression [${}^{+}$] into the algorithm design.
@@ Acknowledgement
This work is supported in part by the grants ONR N00014-19-1-2404, NSF CIF-2007911, CNS-2148212, ECCS-2318441, and AFRL FA8750-20-2-0504. He Wang is also gratefully supported by the Bob Lee Gregory Fellowship at Carnegie Mellon University.
Appendix A Proof of Theorem 1
In this section, we present the convergence analysis of PISCO. For the ease of the analysis, we begin with rewriting the updates of PISCO  in a more compact form. First, by unrolling the local updates (6) in PISCO, we have
@@ Appendix A Proof of Theorem 1
$\boldsymbol{Y}^{k+1}$ $=\left(\boldsymbol{Y}^{k}++\boldsymbol{G}^{k+1}-\boldsymbol{G}^{k+1,T_{o}}+\sum_{t=1}^{T_{o}}\left(\boldsymbol{G}^{k+1,t}-\boldsymbol{G}^{k+1,t-1}\right)\right)\boldsymbol{W}^{k}$
@@ A.1 Useful inequalities
Before proceeding, we first introduce several propositions that will be used in the following analysis.
@@ A.2 Property of gradient tracking
In this section, we first show that $\overline{\boldsymbol{Y}}^{k}=\boldsymbol{Y}^{k}\boldsymbol{J}$ is able to track the average of local stochastic gradients, i.e., $\overline{\boldsymbol{G}}^{k}=\boldsymbol{G}^{k}\boldsymbol{J}=\left(\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{g}_{i}^{k}\right)\boldsymbol{1}_{n}^{\top}$. Define the average model estimate as $\overline{\boldsymbol{x}}^{k}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}\in\mathbb{R}^{d}$ and $\overline{\boldsymbol{X}}^{k}=\boldsymbol{X}^{k}\boldsymbol{J}$.
@@
$\overline{\boldsymbol{Y}}^{k}=\overline{\boldsymbol{G}}^{k},$ (14)
@@
$\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]\leq\frac{3\sigma^{2}}{b}+3L^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]+3n\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}].$ (15)
where $\boldsymbol{\Phi}_{x}^{k}=\boldsymbol{X}^{k}-\overline{\boldsymbol{X}}^{k}$ denotes the consensus error at iteration $k$.
@@
due to the initialization of Algorithm 1 s.t. $\boldsymbol{Y}^{0}=\boldsymbol{G}^{0}$. Moreover,
@@
where $(a)$ is based on $\|\boldsymbol{x}\mathbf{1}_{n}^{\top}\|_{\mathsf{F}}^{2}=n\|\boldsymbol{x}\|_{2}^{2}$, $(b)$ uses (12) in Proposition 1, $(c)$ follows Assumption 2 and 3 and applies Proposition 3. ∎
@@ A.3 Progress improvement on successive iterates and averages
Before continuing, we first show two auxiliary lemmas to bound the progress improvement between the successive iterates and their averages. Similar to $\boldsymbol{\Phi}_{x}^{k}=\boldsymbol{X}^{k}-\overline{\boldsymbol{X}}^{k}$, we also use $\boldsymbol{\Phi}_{y}^{k}=\boldsymbol{Y}^{k}-\overline{\boldsymbol{Y}}^{k}$ to represent the tracking error at the $k$-th iteration. As for the local updates, we define $t$-th local-update consensus error as $\boldsymbol{\Phi}_{x}^{k,t}=\boldsymbol{X}^{k+1,t}-\overline{\boldsymbol{X}}^{k}$ and tracking error as $\boldsymbol{\Phi}_{y}^{k,t}=\boldsymbol{Y}^{k+1,t}-\overline{\boldsymbol{Y}}^{k}$.
@@
$\mathbb{E}\left[\big{\|}\overline{\boldsymbol{X}}^{k+1}-\overline{\boldsymbol{X}}^{k}\big{\|}_{\mathsf{F}}^{2}\right]\leq\frac{3T_{o}\eta^{2}\sigma^{2}}{b}+3T_{o}L^{2}\eta^{2}\sum_{t=0}^{T_{o}}\mathbb{E}\left[\big{\|}\boldsymbol{\Phi}_{x}^{k,t}\big{\|}_{\mathsf{F}}^{2}\right]+3nT_{o}^{2}\eta^{2}\mathbb{E}\left[\big{\|}\nabla f(\overline{\boldsymbol{x}}^{k})\big{\|}_{2}^{2}\right].$ (21)
@@ A.4 Bounding consensus errors and tracking errors
In this subsection, we present three key lemmas for controlling consensus and tracking errors, which will be used to establish the convergence results of PISCO.
We begin with bounding the accumulated consensus errors for local updates in the following lemma.
@@
$\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]$ $\leq(1-p)\lambda^{2}\left[c_{0}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]+2(1+\frac{1}{\delta_{1}})(T_{o}+1)^{2}\eta^{2}\left(32(T_{o}+1)^{2}L^{2}\eta_{l}^{2}+1\right)\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]\right.$
@@
where $c_{0}=1+\delta_{1}+80(1+\frac{1}{\delta_{1}})T_{o}^{2}L^{2}\eta^{2}$. Let $\delta_{1}=\frac{\lambda_{p}}{8}\in(0,1]$. If $\eta\leq\frac{\lambda_{p}}{80(T_{o}+1)L}$, we have
$c_{0}\leq 1+\frac{\lambda_{p}}{4}\quad\text{and}\quad 1+\frac{1}{\delta_{1}}<\frac{10}{\lambda_{p}}.$
@@
Suppose Assumption 1,2 and 3 hold. If $\eta_{l}\leq\frac{1}{8L(T_{o}+1)}$ and $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$,
$\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]$ $\leq(1-p)\left[\frac{1+(1+p)\lambda^{2}}{2}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]+\frac{400\lambda^{2}}{\lambda_{p}}L^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]\right]$
$\quad+\frac{(1-p)\lambda^{2}}{\lambda_{p}}\left[125L^{2}\eta^{2}(T_{o}+1)^{2}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k-1}\|_{\mathsf{F}}^{2}]+180n\sigma^{2}/b\right].$
@@
$\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]$ $\leq(1-p)\lambda^{2}\left[c_{1}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]+36(1+\frac{1}{\delta_{1}})(1+20T_{o}^{2}L^{2}\eta^{2})L^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]+6n(1+\frac{1}{\delta_{1}})\right.$
@@
where $c_{1}=1+\delta_{1}+36(1+\frac{1}{\delta_{1}})(T_{o}+1)^{2}L^{2}\eta^{2}$. If $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$ and $\delta_{1}=\frac{\lambda_{p}}{8}\in(0,1)$,
$c_{1}\leq 1+\frac{\lambda_{p}}{4}\quad\text{and}\quad 1+\frac{1}{\delta_{1}}\leq\frac{10}{\lambda_{p}}.$
@@
Since $\eta_{l}\leq\frac{1}{8L(T_{o}+1)}$ and $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$, we arrive at Lemma 6. ∎

### v4

@@ 5.1 Logistic regression with nonconvex regularization
To investigate communication efficiency of PISCO, we conduct experiments on logistic regression with a nonconvex regularization term [WJZ+19] using the a9a dataset [CL11]. Given the model parameter $\boldsymbol{x}$ and data sample $\boldsymbol{z}=(\boldsymbol{a},y)$, the empirical loss $\ell(\boldsymbol{x};\boldsymbol{z})$ is defined as:
@@
First, we study the influence of the agent-to-server communication probability $p$ on the training and test performance. To this end, we vary the probability $p$ from $\{1,1/10^{0.5},1/10^{0.75},1/10,1/10^{1.25},1/10^{1.5},1/10^{1.75},1/10^{2},0\}$ and present the number of communication rounds of PISCO with different $p$ to achieve $0.05$ training accuracy (i.e., $\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}\leq 0.05$) and $80\%$ test accuracy ($\geq 95\%$ of the peak accuracy within 1000 communication rounds), in Figure 4.
@@
Figure 4: The number of agent-to-agent and agent-to-server communication rounds required to achieve $0.05$ training accuracy (the left panel) and $80\%$ test accuracy (the right panel) for PISCO with $T_{o}=1$ and different $p\in\{0,10^{-2},10^{-1.75},10^{-1.5},10^{-1.25},10^{-1},10^{-0.75},10^{-0.5},1\}$. Here, the blue (red) dotted line represents the number of agent-to-agent (agent-to-server) communication rounds that PISCO with $p=0$ (with $p=1$) requires.
@@
5.2 Neural network training
(b) A disconnected Erdős-Rényi graph with $0.1$ connectivity probability.
@@ 5.2 Neural network training
To verify the robustness of PISCO to diverse topologies, we consider a well-connected Erdős-Rényi topology with a connectivity probability of $0.3$ (corresponding to $\lambda_{w}=0.38$) and a disconnected Erdős-Rényi topology with a connectivity probability of $0.1$ (corresponding to $\lambda_{w}=0$). To simulate the highly data-heterogeneous scenario, we evenly split the sorted MNIST dataset to $n=10$ agents, where agent $i\in[n]$ will receive the training data associated with label $i$. Moreover, we set the batch size $b=100$, the number of local updates $T_{o}=10$ and the agent-to-server communication probability $p\in\{1,1/\sqrt{n},1/n,0\}=\{1,10^{-0.5},10^{-1},0\}$. To reduce the impact of randomness, we run every experiment with 3 different seeds and show the average results in Figure 6.
@@
We also evaluate the performance of PISCO by training a convolutional neural network (CNN) on the unshuffled CIFAR10 dataset [KH+09]. The network architecture includes three sequential CNN modules, each containing two 2D convolutional layers with ReLU activation, followed by max pooling (kernel size 2, stride 2) and dropout (rate 0.2) for regularization. Specifically, in the first module, the initial convolutional layer transforms the input from 3 to 32 channels, and the second convolutional layer maintains 32 channels; the second module follows this pattern, mapping 32 to 64 channels; and the third similarly increases from 64 to 128 channels. After feature extraction, the fully connected layers process the 2048 flattened output, first mapping it to 128 features with ReLU activation and dropout, and then to 10 outputs for classification. We use a ring topology with $n=5$ agent and set the batch size $b=20$ and the number of local updates $T_{o}=4$. To introduce data heterogeneity, we split the sorted CIFAR10 dataset across the 5 agents, so that each agent $i\in[n]$ obtains training data with label $i$ and $i+5$.
In Figure 7, we illustrate the effectiveness of PISCO, in terms of training loss and test accuracy across epochs. We observe that, due to sparse agent-to-agent communication in the ring topology and extremely high data heterogeneity, PISCO with $p=0$ converges more slowly than PISCO with $p>0$. Notably, PISCO with $p=1/\sqrt{5}$ achieves performance comparable to PISCO with $p=1$, demonstrating the efficiency of the heterogeneous communication protocol in reducing costly agent-to-server communications.
Figure 7: The training loss and testing accuracy across epochs with different probabilities $p=1,1/\sqrt{5},0.2,0$ and the number of local updates $T_{o}=4$, over a ring topology for CNN training on the sorted CIFAR10 dataset.
@@ 6 Conclusions
In this paper, we develop a communication-efficient algorithm called PISCO for solving federated nonconvex optimization over semi-decentralized networks, which enjoys the linear speedup of local updates and addresses data dissimilarity without any additional assumptions. By leveraging the heterogeneous communication model, PISCO largely reduces communication overheads in terms of the network dependency with a few agent-to-server communication rounds, particularly evident in poorly-connected networks. Both theoretical guarantees and empirical experiments underscore PISCO’s outstanding communication efficiency and robustness to data heterogeneity and various network topologies.
In the future, it will be of interest to incorporate variance reduction techniques [LCCC20, BBN18] into the algorithm design, apply communication compression [ZLL+22] to further reduce the per-round communication costs, and enable varying agent-to-server communication probabilities [SSY+24], allowing for personalized and heterogeneous communication strategies for each agent.
@@ Acknowledgement
This work is supported in part by the grants ONR N00014-19-1-2404, NSF CIF-2007911, CNS-2148212, ECCS-2318441, AFRL FA8750-20-2-0504, and is supported in part by funds from federal agency and industry partners as specified in the Resilient & Intelligent NextG Systems (RINGS) program. He Wang is also gratefully supported by the Bob Lee Gregory Fellowship at Carnegie Mellon University.
Appendix A Technical Lemmas
This section establishes several critical lemmas which will be used in the proof of Theorem 1. For the ease of the analysis, we begin with rewriting the updates of PISCO  in a more compact form. First, by unrolling the local updates (3) in PISCO, we have
@@ Appendix A Technical Lemmas
$\boldsymbol{Y}^{k+1}$ $=\left(\boldsymbol{Y}^{k}+\boldsymbol{G}^{k+1}-\boldsymbol{G}^{k+1,T_{o}}+\sum_{t=1}^{T_{o}}\left(\boldsymbol{G}^{k+1,t}-\boldsymbol{G}^{k+1,t-1}\right)\right)\boldsymbol{W}^{k}$
@@ A.1 Useful inequalities
Before proceeding, we introduce the following propositions which will be useful in the analysis.
@@ A.2 Property of gradient tracking
In this section, the following lemma shows that $\overline{\boldsymbol{Y}}^{k}=\boldsymbol{Y}^{k}\boldsymbol{J}$ is able to track the average of local stochastic gradients, i.e., $\overline{\boldsymbol{G}}^{k}=\boldsymbol{G}^{k}\boldsymbol{J}=\left(\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{g}_{i}^{k}\right)\boldsymbol{1}_{n}^{\top}$. We define the average model estimate as $\overline{\boldsymbol{x}}^{k}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}^{k}\in\mathbb{R}^{d}$ and $\overline{\boldsymbol{X}}^{k}=\boldsymbol{X}^{k}\boldsymbol{J}$.
@@
$\overline{\boldsymbol{Y}}^{k}=\overline{\boldsymbol{G}}^{k}.$
@@
$\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k}\|_{\mathsf{F}}^{2}]\leq\frac{3\sigma^{2}}{b}+3L^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]+3n\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}],$ (12)
where $\boldsymbol{\Phi}_{x}^{k}=\boldsymbol{X}^{k}-\overline{\boldsymbol{X}}^{k}$ is the consensus error at iteration $k$.
@@
due to the initialization of Algorithm 1 s.t. $\boldsymbol{Y}^{0}=\boldsymbol{G}^{0}$. Moreover, we define the distributed objective as $F(\boldsymbol{X})=\sum_{i=1}^{n}f_{i}(\boldsymbol{x}_{i})$, where the distributed gradient is denoted by $\nabla F(\boldsymbol{X})=\left[\nabla f_{1}(\boldsymbol{x}_{1}),\nabla f_{2}(\boldsymbol{x}_{2}),\ldots,\nabla f_{n}(\boldsymbol{x}_{n})\right]\in\mathbb{R}^{d\times n}.$ Then, we have
@@
where $(a)$ is based on $\|\boldsymbol{x}\mathbf{1}_{n}^{\top}\|_{\mathsf{F}}^{2}=n\|\boldsymbol{x}\|_{2}^{2}$, $(b)$ uses (10) in Proposition 1, $(c)$ follows Assumption 2 and 3 and applies Proposition 3.
@@ A.3 Progress improvement on successive iterates and averages
The following two auxiliary lemmas bound the progress improvement between the successive iterates and their averages. Similar to $\boldsymbol{\Phi}_{x}^{k}=\boldsymbol{X}^{k}-\overline{\boldsymbol{X}}^{k}$, we also use $\boldsymbol{\Phi}_{y}^{k}=\boldsymbol{Y}^{k}-\overline{\boldsymbol{Y}}^{k}$ to represent the tracking error at the $k$-th iteration. As for the local updates, we define $t$-th local-update consensus error as $\boldsymbol{\Phi}_{x}^{k,t}=\boldsymbol{X}^{k+1,t}-\overline{\boldsymbol{X}}^{k}$ and tracking error as $\boldsymbol{\Phi}_{y}^{k,t}=\boldsymbol{Y}^{k+1,t}-\overline{\boldsymbol{Y}}^{k}$.
@@
$\mathbb{E}[\|\boldsymbol{X}^{k}-\boldsymbol{X}^{k-1}\|_{\mathsf{F}}^{2}]$ $\leq 12\left(1+2T_{o}^{2}L^{2}\eta^{2}\right)\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]+6(1-p)\lambda^{2}(T_{o}+1)^{2}\eta^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]+\frac{48nT_{o}^{2}\eta^{2}\sigma^{2}}{b}$
$+24T_{o}L^{2}\eta^{2}\sum_{t=1}^{T_{o}}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1,t}\|_{\mathsf{F}}^{2}]+3(T_{o}+1)^{2}\eta^{2}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k-1}\|_{\mathsf{F}}^{2}].$
@@
$\mathbb{E}\!\left[\big{\|}\overline{\boldsymbol{X}}^{k+1}-\overline{\boldsymbol{X}}^{k}\big{\|}_{\mathsf{F}}^{2}\right]$ $\leq\frac{3T_{o}\eta^{2}\sigma^{2}}{b}\!+\!3T_{o}L^{2}\eta^{2}\!\sum_{t=0}^{T_{o}}\mathbb{E}\left[\big{\|}\boldsymbol{\Phi}_{x}^{k,t}\big{\|}_{\mathsf{F}}^{2}\right]+3nT_{o}^{2}\eta^{2}\mathbb{E}\left[\big{\|}\nabla f(\overline{\boldsymbol{x}}^{k})\big{\|}_{2}^{2}\right].$
@@ A.4 Bounding consensus errors and tracking errors
Next, we present the following lemma for bounding the accumulated consensus errors for local updates, in order to control the consensus and tracking errors at every iteration.
@@
$\quad+\frac{240\lambda^{2}L^{2}(T_{o}+1)^{4}\eta^{2}\eta_{l}^{2}}{\lambda_{p}}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k-1}\|_{\mathsf{F}}^{2}]\left.+\frac{320\lambda^{2}n(T_{o}+1)^{2}\eta^{2}}{\lambda_{p}}\frac{\sigma^{2}}{b}\right].$
@@
$\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k}\|_{\mathsf{F}}^{2}]$ $\leq(1-p)\lambda^{2}\left[\theta_{0}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]+2(1+\frac{1}{\delta_{1}})(T_{o}+1)^{2}\eta^{2}\left(32(T_{o}+1)^{2}L^{2}\eta_{l}^{2}+1\right)\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]\right.$
@@
where $\theta_{0}=1+\delta_{1}+80(1+\frac{1}{\delta_{1}})T_{o}^{2}L^{2}\eta^{2}$. Let $\delta_{1}=\frac{\lambda_{p}}{8}\in(0,1]$. If $\eta\leq\frac{\lambda_{p}}{80(T_{o}+1)L}$, we have
$\theta_{0}\leq 1+\frac{\lambda_{p}}{4}\quad\text{and}\quad 1+\frac{1}{\delta_{1}}<\frac{10}{\lambda_{p}}.$
@@
$\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]$ $\leq(1-p)\lambda^{2}\!\left[\frac{1+(1+p)}{2}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]\!+\!\frac{400}{\lambda_{p}}L^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]\right]$
$\quad+\frac{(1-p)\lambda^{2}}{\lambda_{p}}\left[125L^{2}\eta^{2}(T_{o}+1)^{2}\mathbb{E}[\|\overline{\boldsymbol{Y}}^{k-1}\|_{\mathsf{F}}^{2}]+180n\frac{\sigma^{2}}{b}\right].$
@@
$\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k}\|_{\mathsf{F}}^{2}]$ $\leq(1-p)\lambda^{2}\left[\theta_{1}\mathbb{E}[\|\boldsymbol{\Phi}_{y}^{k-1}\|_{\mathsf{F}}^{2}]+36(1+\frac{1}{\delta_{1}})(1+20T_{o}^{2}L^{2}\eta^{2})L^{2}\mathbb{E}[\|\boldsymbol{\Phi}_{x}^{k-1}\|_{\mathsf{F}}^{2}]+6n(1+\frac{1}{\delta_{1}})\right.$
@@
where $\theta_{1}=1+\delta_{1}+36(1+\frac{1}{\delta_{1}})(T_{o}+1)^{2}L^{2}\eta^{2}$. If $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$ and $\delta_{1}=\frac{\lambda_{p}}{8}\in(0,1)$,
$\theta_{1}\leq 1+\frac{\lambda_{p}}{4}\quad\text{and}\quad 1+\frac{1}{\delta_{1}}\leq\frac{10}{\lambda_{p}}.$
@@
Since $\eta_{l}\leq\frac{1}{8L(T_{o}+1)}$ and $\eta\leq\frac{\lambda_{p}}{80L(T_{o}+1)}$, we arrive at Lemma 6.